OpenStudy (anonymous):

Pleas help!!!!!!!!!!!!!! Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (anonymous):

@ParthKohli you're a genius ;D you help hehe

OpenStudy (anonymous):

i dnt get it

OpenStudy (anonymous):

Oh, I tagged someone so they will come view your problem when you do "@" and someone's username it sends them a notification and a link to where they were tagged. So, hopefully @ParthKohli will come help you. He's a mathematical genius. :)

OpenStudy (anonymous):

o lol ok

OpenStudy (anonymous):

i would use substitution... 4x +3y = -13x + y therefore sub the second equation into the first to get 4x - 39x = 3, therefore -35x = 3 therfore x = -3/35. now sub this value into the very first equation to find y and there you go ... both equations or pronumerals solved.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

do you get it now?... you basically sub one equation into another to find out one vale and once you knowone value you only have unknown left which can be simply solved useng basic algebra.

OpenStudy (cwrw238):

elimination would also be quite easy to use multiply 2nd equation by -3 and then add will eliminate y

OpenStudy (anonymous):

I don't know what this is called but I solved it using sbstraction sub 4x+3y=-1 from 3x+y=3 ans is x=2